# Tianjun Gu1 & Linfeng Li1 & Xuhong Wang2 & Chenghua Gong1 & Jingyu Gong1 & Zhizhong Zhang1 & Yuan Xie1,3 & Lizhuang Ma1 & Xin Tan1,2

### v3

@@ Appendix E Experimental Setup Details
The maximal navigation steps per episode are set to 40. The agent’s body has a radius of $0.17m$ and a height of $1.5m$. Its RGB-D sensors are positioned at $1.5m$ height with a $-0.45$ radian downward tilt and provide a $131^{\circ}$ Field of View (FoV). For rotation, the agent selects an angular displacement corresponding to one of 60 discrete bins that uniformly discretize the $360^{\circ}$ range. Success requires stopping within $d_{\text{success}}=0.3$m of the target object and visually confirming it. Success requires stopping within $d_{\text{success}}=0.3m$ of the target object and visually confirming it. Our DORAEMON framework primarily utilizes Gemini-1.5-pro as the VLM and CLIP ViT-B/32 for semantic embeddings, with caching implemented for efficiency. Key hyperparameters include: topological map connection distance $\delta_{\text{connect}}=1.0$m, node update interval $S_{\text{update}}=3$ steps, $L_{1}$ hierarchical clustering weight $w=0.4$, AORI grid resolution $\delta_{\text{grid}}=0.1$m, minimum obstacle clearance $d_{\text{min\_obs}}=0.5$m, and various stuck detection thresholds (e.g., path inefficiency $\eta_{\text{path}}<0.25$, small area coverage $\delta_{\text{area\_gain}}<0.35$m2, high rotation/translation ratio $\rho_{\text{rot/trans}}>2.0$ for short paths when $\|\text{path}\|<0.5$m) and a precision movement factor $\gamma_{\text{step}}=0.1$.
@@ F.1 Level $L_{3}$: Observation Anchoring
[…]
@@ Appendix H Chain-of-Thought Prompt
Our Policy-VLM leverages a structured Chain-of-Thought (CoT) prompt to guide the decision-making process. The complete prompt is provided below:
@@ Appendix I Detailed Description of Baseline
To assess the performance of DORAEMON, we compare it with 16 recent baselines for (zero-shot) object-goal navigation. Summaries are given below.
ProcTHOR [12]: A procedurally–generated 10K-scene suite for large-scale Embodied AI.
ProcTHOR_ZS [12]: ProcTHOR_ZS trains in ProcTHOR and evaluates zero-shot on unseen iTHOR/RoboTHOR scenes to test cross-domain generalisation.
SemEXP [9]: Builds an online semantic map and uses goal-oriented exploration to locate the target object efficiently, achieving state-of-the-art results in Habitat ObjectNav 2020.
Habitat-Web [31]: Collects large-scale human demonstrations via a browser interface and leverages behaviour cloning to learn object-search strategies.
PONI [28]: Learns a potential-field predictor from static supervision, enabling interaction-free training while preserving high navigation success.
ZSON [22]: Encodes multimodal goal embeddings (text + images) to achieve zero-shot navigation towards previously unseen object categories.
PSL [35]: Prioritised Semantic Learning selects informative targets during training and uses semantic expansion at inference for zero-shot instance navigation.
Pixel-Nav [5]: Introduces pixel-guided navigation skills that bridge foundation models and ObjectNav, relying solely on RGB inputs.
SGM [47]: “Imagine Before Go” constructs a self-supervised generative map to predict unseen areas and improve exploration efficiency.
ImagineNav [48]: Prompts vision–language models to imagine future observations, guiding the agent toward information-rich viewpoints.
CoW [13]: Establishes the “Cows on Pasture” benchmark for language-driven zero-shot ObjectNav and releases baseline policies without in-domain training.
ESC [51]: Employs soft commonsense constraints derived from language models to bias exploration, markedly improving zero-shot success over CoW.
@@
VLFM [42]: Combines VLM goal-localisation with frontier-based exploration, removing the need for reinforcement learning or task-specific fine-tuning.
VoroNav [39]: Simplifies the search space via Voronoi partitions and pairs this with LLM-driven semantic planning for improved zero-shot performance.
TopV-Nav [50]: Lets a multimodal LLM perform spatial reasoning directly on top-view maps, with adaptive visual prompts for global–local coordination.
SG-Nav [41]: Online builds a 3D scene graph and uses hierarchical Chain--of--Thought prompting so an LLM can infer probable target locations.

### v4

@@ Appendix E Experimental Setup Details
The maximal navigation steps per episode are set to 40. The agent’s body has a radius of $0.17m$ and a height of $1.5m$. Its RGB-D sensors are positioned at $1.5m$ height with a 0.45-0.45 radian downward tilt and provide a $131^{\circ}$ Field of View (FoV). For rotation, the agent selects an angular displacement corresponding to one of 60 discrete bins that uniformly discretize the $360^{\circ}$ range. Success requires stopping within $d_{\text{success}}=0.3$m of the target object and visually confirming it. Success requires stopping within $d_{\text{success}}=0.3m$ of the target object and visually confirming it. Our DORAEMON framework primarily utilizes Gemini-1.5-pro as the VLM and CLIP ViT-B/32 for semantic embeddings, with caching implemented for efficiency. Key hyperparameters include: topological map connection distance $\delta_{\text{connect}}=1.0$m, node update interval $S_{\text{update}}=3$ steps, $L_{1}$ hierarchical clustering weight $w=0.4$, AORI grid resolution $\delta_{\text{grid}}=0.1$m, minimum obstacle clearance $d_{\text{min\_obs}}=0.5$m, and various stuck detection thresholds (e.g., path inefficiency $\eta_{\text{path}}<0.25$, small area coverage $\delta_{\text{area\_gain}}<0.35$m2, high rotation/translation ratio $\rho_{\text{rot/trans}}>2.0$ for short paths when $\|\text{path}\|<0.5$m) and a precision movement factor $\gamma_{\text{step}}=0.1$.
@@ F.3 Level $L_{1}$: Room Formation ($L_{2}$$L_{1}$)
      $\mathcal{F}_{s}=\{\mathcal{A}_{s,f}|f=\text{MapToRoomFunction}(\texttt{area\_type})\}.$ (20)
@@ Appendix H Chain-of-Thought Prompt
Our Exec-VLM leverages a structured Chain-of-Thought (CoT) prompt to guide the decision-making process. The complete prompt is provided below:
@@ Appendix I Detailed Description of Baseline
To assess the performance of DORAEMON, we compare it with 16 recent baselines for (zero‑shot) object‑goal navigation. Summaries are given below.
ProcTHOR [12]: A procedurally–generated 10K‑scene suite for large‑scale Embodied AI.
ProcTHOR_ZS [12]: ProcTHOR_ZS trains in ProcTHOR and evaluates zero‑shot on unseen iTHOR/RoboTHOR scenes to test cross‑domain generalisation.
SemEXP [9]: Builds an online semantic map and uses goal‑oriented exploration to locate the target object efficiently, achieving state‑of‑the‑art results in Habitat ObjectNav 2020.
Habitat‑Web [30]: Collects large‑scale human demonstrations via a browser interface and leverages behaviour cloning to learn object‑search strategies.
PONI [27]: Learns a potential‑field predictor from static supervision, enabling interaction‑free training while preserving high navigation success.
ZSON [21]: Encodes multimodal goal embeddings (text + images) to achieve zero‑shot navigation towards previously unseen object categories.
PSL [34]: Prioritised Semantic Learning selects informative targets during training and uses semantic expansion at inference for zero‑shot instance navigation.
Pixel‑Nav [5]: Introduces pixel‑guided navigation skills that bridge foundation models and ObjectNav, relying solely on RGB inputs.
SGM [45]: “Imagine Before Go” constructs a self‑supervised generative map to predict unseen areas and improve exploration efficiency.
ImagineNav [46]: Prompts vision–language models to imagine future observations, guiding the agent toward information‑rich viewpoints.
CoW [13]: Establishes the “Cows on Pasture” benchmark for language‑driven zero‑shot ObjectNav and releases baseline policies without in‑domain training.
ESC [49]: Employs soft commonsense constraints derived from language models to bias exploration, markedly improving zero‑shot success over CoW.
@@
VLFM [41]: Combines VLM goal‑localisation with frontier‑based exploration, removing the need for reinforcement learning or task‑specific fine‑tuning.
VoroNav [37]: Simplifies the search space via Voronoi partitions and pairs this with LLM‑driven semantic planning for improved zero‑shot performance.
TopV‑Nav [48]: Lets a multimodal LLM perform spatial reasoning directly on top‑view maps, with adaptive visual prompts for global–local coordination.
SG‑Nav [39]: Online builds a 3D scene graph and uses hierarchical Chain‑-of-‑Thought prompting so an LLM can infer probable target locations.